Bill can row 3 mph in still water. It takes him 3 hours 36 minutes to go 3 miles upstream and return. Find the speed of the current.
I've gone over and over trying this problem different ways. I've converted the time into 216 minutes and still can't figure out what I'm doing wrong.

Answer by Alan3354(69443)

Bill can row 3 mph in still water. It takes him 3 hours 36 minutes to go 3 miles upstream and return. Find the speed of the current.
---------------------
current speed = c
d = rt = 3 miles
t = 3/r
3.6 = 3/(3+c) + 3/(3-c)
3.6(3+c)*(3-c) = 3(3-c) + 3(3+c)
1.2(9 - c^2) = 3-c + 3+c = 6
9 - c^2 = 5
c^2 = 4
c = 2 mph
---------
It takes 3 hour upstream at 1 mph ground speed, and
0.6 hours downdstream at 5 mph.